# Conditional Feature Importance revisited: Double Robustness, Efficiency and Inference

### v4

@@ Abstract
Conditional Feature Importance (CFI) was introduced long ago to account for the relationship between the studied feature and the rest of the input. However, CFI has not yet been studied from a theoretical perspective because the conditional sampling step has generally been overlooked. In this article, we demonstrate that the recent Conditional Permutation Importance (CPI) is indeed a valid implementation of this concept. Under the conditional null hypothesis, we then establish a double robustness property that can be leveraged for variable selection: with either a valid model or a valid conditional sampler, the method correctly identifies null coordinates.
@@ 1 Introduction
Modern machine learning models are capable of making accurate predictions, so using these models can help characterize the dependency structure between variables and thus provide valuable insights for further research. For example, if a model can accurately predict a disease based on a set of genes, identifying the important predictive variables can guide specialists on which genes to study further to combat the disease.
@@
There is usually a trade-off between model transparency and model complexity. Indeed, simple methods such as linear regression are easily interpretable, since each covariate’s importance can be directly assessed via its estimated coefficient. However, using a model that poorly fits the underlying distribution may yield misleading results (Molnar et al.,, 2021).
For this reason, there is a need to rely on model-agnostic approaches based on flexible enough models that truly provide information about the data distribution. In this way, one can adapt to complex situations by capturing non-linear dependencies, recovering the truly important variables in highly correlated settings.
@@
The two main model-agnostic approaches to measure the importance of a covariate are perturbation and removal approaches (Covert et al.,, 2021). Both of them aim to disable the information given by the covariate.
@@
In perturbation/permutation approaches, such as Permutation Feature Importance (PFI, Mi et al., (2021)) and Conditional Permutation Importance (CPI, Chamma et al., 2024a ), the information is disabled by computing the loss when permuting the $j$-th coordinate marginally or conditionally.
@@
The most used removal-based approach is Leave One Covariate Out (LOCO), where the importance of a given covariate is estimated by evaluating the performance of a model re-trained without that covariate (Lei et al.,, 2018; Williamson et al.,, 2021).
@@
LOCO is essentially a direct plug-in estimate of the predictive power of the model compared to the predictive power of the restricted model (Williamson et al.,, 2023). As such, it effectively estimates the Total Sobol Index (TSI), an importance index originating from Sensitivity Analysis (Homma and Saltelli,, 1996). TSI aims to assess the predictive capacity of a coordinate given the others, making it a conditional approach.
@@
Due to extrapolation issues with PFI (Hooker et al.,, 2021), there has been a shift towards conditional approaches. These include Conditional Variable Importance—also called Conditional Feature Importance (CFI) in Ewald et al., (2024)—and Conditional Model Reliance (CMR), defined as a ratio rather than a difference in Fisher et al., (2019). However, the conditional sampling step is often overlooked; a Gaussian distribution or copula is commonly used but performs poorly in practice due to covariance estimation issues (Blain et al.,, 2025). Our first contribution is to show that the conditional permutation used in CPI (and CMR) has theoretical grounding, making it a valid CFI approach.
@@
These conditional methods have shown strong empirical performance (see Chamma et al., 2024a ; Chamma et al., 2024b ; Paillard et al., (2025)). We explain this by a double robustness property, typically sought in causal inference, which we introduce here for variable importance.
@@
Variable selection seeks the smallest set of predictive variables—those not independent of the output given the others. Even in highly correlated settings, assuming that each covariate is not directly a function of the others, this set is well-defined and unique. This assumption is standard in controlled variable selection (Candès et al.,, 2018) and feature importance (Verdinelli and Wasserman,, 2024).
@@
Permutation-based approaches do not provide a proper quantification of variable importance. In particular, Bénard et al., (2022) have shown that PFI does not converge to TSI. Yet, the theoretical quantity targeted by CPI has not been studied. In this article, we fill this gap and show that, with a simple adjustment, the method can be adapted to estimate TSI. We call this extension Sobol-CPI, an unbiased generalization of CPI, for which we establish nonparametric efficiency and study inference to provide statistical guarantees.
@@
  • a theoretical framework under which the conditional sampling step of CPI is valid, which shows that CPI is a Conditional Feature Importance.
  • a double robustness property: to detect a null covariate, the accuracy of only one of the two estimates used in its computation is sufficient. This explains the strong performance of CPI in variable selection. By contrast, LOCO requires both full and restricted models to be accurate. In a simple linear setting, CPI exhibits quadratic bias decay, whereas LOCO decays only linearly.
  • Sobol-CPI, a new estimator of TSI based on a permutation approach, bridging the gap with removal methods. We prove its asymptotic efficiency and provide a consistent (power tending to one) and valid (type-I error control) test.
  • numerical experiments to illustrate these findings.
Proofs and additional experiments are in the appendix.
@@ 2.1 Setting
We observe $\{(x_{i},y_{i})\}_{i=1,\ldots,n_{\mathrm{test}},\ldots,n_{\mathrm{test}}+n_{\mathrm{train}}}$ a test and train set sampled i.i.d. from $P_{0}$, a distribution that belongs to a class of distributions $\mathcal{M}$, where $X\in\mathcal{X}\subset\mathbb{R}^{p}$ and $y\in\mathcal{Y}\subset\mathbb{R}$. We are interested in assessing the importance of a coordinate $j\in\{1,\ldots,p\}$. We denote by $X^{j}$ the $j$-th column of $X$ and by $X^{-j}$ the vector $X$ with the $j$-th coordinate excluded. We consider a space of functions $\mathcal{F}$ with a norm $\|\cdot\|_{\mathcal{F}}$. The notation is summarized in the glossary (Appendix A).
Similarly to the Knockoffs framework (Barber and Candès,, 2015; Candès et al.,, 2018), we want to avoid making strong assumptions about the relationship between inputs and outputs, as it can be complex. However, we assume that the relationship among the input covariates is simple, typically because they originate from the same generative process (e.g. a measurement device). For this reason, it can be much more efficient and accurate to study the relationship between the $j$-th covariate, $X^{j}$, given the rest, $X^{-j}$, rather than directly regressing $y$ on $X^{-j}$. This approach is also advantageous in settings where labeling data is expensive, as there are typically more available samples for the input pair $(X^{-j},X^{j})$ than for the pair $(X^{-j},y)$. However, empirical robustness to this assumption is demonstrated in Paillard et al., (2025).
We denote by $m(X):=\mathbb{E}[y\mid X]$ (resp. $m_{-j}(X^{-j}):=\mathbb{E}[y\mid X^{-j}]$) the conditional expectation of the input $y$ given $X$ (resp. $X^{-j}$), by $\widehat{m}$ (resp. $\widehat{m}_{-j}$) its estimation and by $\widehat{m}_{n}$ to make explicit the dependence on the training sample size. Similarly, we denote by $\nu_{-j}(X^{-j}):=\mathbb{E}[X^{j}\mid X^{-j}]$ and $\widehat{\nu}_{-j}$ its estimate. We denote by $\ell:\widehat{\mathcal{Y}}\times\mathcal{Y}\to\mathbb{R}$ the loss used to compare a prediction $\widehat{m}(X)\in\widehat{\mathcal{Y}}\subset\mathbb{R}$ with the true output $y\in\mathcal{Y}$. These estimations are the usual objective of machine learning models. This is obvious for the mean squared error (MSE), but for many other losses ($R^{2}$, deviance, classification accuracy, and the area under the ROC curve, see Williamson et al., (2023)) the minimizer is also a function of this conditional expectation.
@@
We define this quantity as the (Generalized) Total Sobol Index because, when the loss is the quadratic loss, it corresponds to the unnormalized TSI (Homma and Saltelli,, 1996). It represents the loss decay when the $j$-th covariate is not used. It has also been referred to as the Generalized ANOVA (Williamson et al.,, 2021) and is considered a standard importance measure (Lei et al.,, 2018; Rinaldo et al.,, 2019; Hooker et al.,, 2021; Bénard et al.,, 2022; Williamson et al.,, 2023). A large decay indicates that the $j$-th covariate is useful, whereas a small decay indicates that the covariate lacks predictive power when the remaining covariates are used. Bénard et al., (2022) considered it as the best quantity for support recovery.
@@ 2.2 Related work
Nonparametric theory (Kennedy,, 2023) shows that a simple plug-in estimate does not achieve optimal convergence, typically requiring a one-step correction. However, Williamson et al., (2021) demonstrated that this correction is unnecessary with the quadratic loss, and Williamson et al., (2023) extended this to other losses under certain regularity conditions. This provides valid confidence intervals. Similar results will be established for our method in Section 4.1.
A major practical limitation of LOCO is that it requires retraining a model, $\widehat{m}_{-j}$, for each coordinate $j$. This is not only computationally intensive, but it also introduces optimization errors that do not compensate as desired, as discussed in Section 3.2.
A first naive permutation-based approach consists in comparing the performance of the estimate on a test set where the $j$-th column is shuffled: the formal definition of PFI is given in Appendix B. Even though it avoids refitting and Mi et al., (2021) report good performance, it suffers from extrapolation bias, as predictions may fall in low-density regions where $\widehat{m}$ was not trained (Hooker et al.,, 2021). It also fails to control type-I error with highly correlated covariates (Chamma et al., 2024a, ). Moreover, Bénard et al., (2022) showed that it does not target an interpretable quantity: it decomposes into three components, only the first being desirable ($\psi_{\mathrm{TSI}}$). The second, the unnormalized marginal total Sobol index, can mislead under high correlation, and the third grows with covariate dependence.
To address this issue, Strobl et al., (2008) proposed conditionally permuting the $j$-th coordinate with respect to the others. In this way, only the information exclusively given by $X^{j}$ is shuffled, while the relationship with the rest of the coordinates is preserved. As a result, one can make predictions based on the new conditional sample, denoted by $\widetilde{x}^{\prime(j)}_{i}$, where the apostrophe indicates that it was estimated. More formally, Chamma et al., 2024a proposed:
@@
CPI requires training a separate regressor $\widehat{\nu}_{-j}$ for each covariate, which may seem counterintuitive since permutation approaches aim to avoid fitting $\widehat{m}_{-j}$. However, as discussed in Section 2.1, while the relationship between $y$ and $X$ may be complex, between $X^{j}$ and $X^{-j}$ is assumed to be simple. For this reason, unlike with LOCO, we use a simple model to estimate $\nu_{-j}$.
The goal of CPI is to sample $\widetilde{X}^{(j)}$ such that $\widetilde{X}^{(j)}\sim X,\widetilde{X}^{(j)-j}=X^{-j}$ and $\widetilde{X}^{(j)j}\perp\!\!\!\perp y\mid X^{-j}$. However, Chamma et al., 2024a did not discuss the assumptions under which this method samples from the target distribution; we address this in Section 3.1. Furthermore, CPI does not estimate a known theoretical quantity, so in Section 4 we propose a modification to target TSI.
Finally, in Chamma et al., 2024a , a type-I error control is proposed using asymptotic normality. However, it does not address the fact that, under the conditional null hypothesis, similarly to what happens with LOCO, the influence function vanishes. Consequently, there is a need to correct the estimated variance to ensure type-I error control (Williamson et al.,, 2023; Dai et al.,, 2024; Verdinelli and Wasserman,, 2024). In Section 4.3, we address this issue.
@@ 3.1 Validity of conditional sampling
We observe that there is no need to preserve the same conditional sampling used for CPI in Chamma et al., 2024a . Normalizing flows (Papamakarios et al.,, 2021) are theoretically sound, but impractical due to the large number of training samples they require. Other alternatives are domain-specific generative models (Sesia et al.,, 2020). Since the conditional sampling used in CPI works in practice and makes the procedure computationally efficient thanks to the use of fast and simple models to estimate $\nu_{-j}$, we continue with the same conditional step, which has also been used in Hooker et al., (2021) and Blain et al., (2025), for example. Let us denote the $i$-th observation in which only the $j$-th coordinate has been conditionally permuted using the residual from the $k$-th observation by:
$\widetilde{x}^{{}^{\prime}(j),l}_{i,k}=\begin{cases}x^{l}_{i}&\text{if }l\neq j\\ \widehat{\nu}_{-j}(x_{i}^{-j})+\left[x_{k}^{j}-\widehat{\nu}_{-j}(x_{k}^{-j})\right]&\text{if }l=j\end{cases}$ (1)
@@
For each $j\in\{1,\ldots,p\}$, there exists a function $\nu_{-j}$ such that $X^{j}=\nu_{-j}(X^{-j})+\epsilon_{j}$ with $\epsilon_{j}\perp\!\!\!\perp X^{-j}$ and $\mathbb{E}\left[\epsilon_{j}\right]=0$.
@@
For instance, the Random Forest is consistent under mild assumptions (Scornet et al., (2015)). If we assume, for instance, that $X$ is Gaussian, a simple linear model is consistent. To deal with high-dimensionality, we generally use a Lasso (Tibshirani, (1996)).
@@
Proposition 3.3 implies that the CPI sampler is asymptotically drawing from the desired conditional distribution. Therefore, CPI qualifies as a Conditional Feature Importance (Hooker et al., (2021)). In the sequel, we refer primarily to CPI due to its feasibility. However, note that the double robustness result from Theorem 3.6, as well as the use of Sobol-CPI to estimate TSI, are completely general and hold for any conditional sampler. Consequently, these results apply to any CFI.

### v5

@@ Abstract
Conditional Feature Importance (CFI; Strobl et al. (2008)) is a classical variable importance measure that accounts for the relationship between the studied feature and the others. However, CFI has not yet been studied from a theoretical perspective because the conditional sampling step has generally been considered a purely practical problem. In this article, we demonstrate that the recent Conditional Permutation Importance (CPI) is indeed a valid implementation of this concept. Under the conditional null hypothesis, we then establish a double robustness property that can be used for variable selection. With either a valid model or a valid conditional sampler, the method correctly identifies null coordinates.
@@
Variable Importance, Conditional Feature Importance, Controlled Variable Selection
@@ 1 Introduction
Modern machine learning models can make accurate predictions, so using them can help characterize the dependency structure between variables and thus provide valuable insights for further research. For example, if a model can accurately predict a disease based on a set of genes, identifying the important predictive variables can guide specialists on which genes to study further to combat the disease.
@@
There is usually a trade-off between model transparency and complexity. Indeed, simple methods such as linear regression are easily interpretable, since each covariate’s importance can be directly assessed via its estimated coefficient. However, using a model that poorly fits the underlying distribution may yield misleading results (Molnar et al., 2021). Therefore, one needs to rely on model-agnostic approaches based on flexible models that give information about the data distribution. In this way, one can adapt to complex situations by capturing non-linear dependencies, recovering the truly important variables in correlated settings.
@@
The two main model-agnostic approaches to measure the importance of a covariate are perturbation and removal approaches (Covert et al., 2021). Both of them aim to disable the information given by the covariate.
@@
In perturbation/permutation approaches, such as Permutation Feature Importance (PFI, Mi et al. (2021)) and Conditional Permutation Importance (CPI, Chamma et al. (2024a)), the information is disabled by computing the loss when permuting the coordinate marginally or conditionally.
@@
The most used removal method is Leave One Covariate Out (LOCO), where the importance of a covariate is estimated by evaluating the performance of a model re-trained without that covariate (Lei et al., 2018; Williamson et al., 2021).
@@
LOCO is essentially a direct plug-in estimate of the predictive power of the model compared to the predictive power of the restricted model (Williamson et al., 2023). As such, it effectively estimates the Total Sobol Index (TSI), an importance index originating from Sensitivity Analysis (Homma and Saltelli, 1996). TSI aims to assess the predictive capacity/ explained variance of a coordinate given the others.
@@
Due to extrapolation issues with PFI (Hooker et al., 2021), there has been a shift towards conditional approaches. These include Conditional Feature Importance (CFI) (Strobl et al., 2008) and Conditional Model Reliance (CMR; Fisher et al. (2019)), defined as a ratio rather than a difference. However, the conditional sampling step is often overlooked; a Gaussian distribution or copula is commonly used but performs poorly in practice due to covariance estimation issues (Blain et al., 2025). Our first contribution is to show that the conditional permutation used in CPI (and CMR) has theoretical grounding, making it an instance of CFI.
@@
These conditional methods have shown strong empirical performance (Chamma et al. (2024a, b); Paillard et al. (2025)). We explain this by a double robustness property, typically sought in causal inference, which we introduce here for variable importance.
@@
Variable selection seeks the smallest set of predictive variables—those not independent of the output given the others. Assuming that each covariate is not directly a function of the others, this set is well-defined and unique. This assumption is standard in variable selection (Candès et al., 2018) and feature importance (Verdinelli and Wasserman, 2024).
@@
Permutation-based approaches do not provide a proper quantification of variable importance. In particular, Bénard et al. (2022) have shown that PFI does not converge to TSI. Yet, the theoretical quantity targeted by CPI has not been studied. In this article, we fill this gap and show that, with a simple adjustment, the method can be adapted to estimate TSI. We call this extension Sobol-CPI, an unbiased generalization of CPI, for which we establish nonparametric efficiency and study inference to provide statistical guarantees.
@@
  • Provide conditions under which the CPI conditional sampler is valid, showing that CPI constitutes a CFI measure.
  • a double robustness property: to detect a null covariate, only one of the two estimates used in CFI needs to be accurate. This explains the empirical performance of CPI. By contrast, LOCO requires both models to be accurate. In linear settings, CPI’s bias decays quadratically, while LOCO’s decays linearly.
  • Sobol-CPI, a new estimator of TSI based on a permutation approach, bridging the gap with removal methods. We prove its asymptotic efficiency.
  • a consistent and valid test using variance correction, and a nonparametric test with finite-sample guarantees based on the sampler’s accuracy.
  • extensive numerical experiments to illustrate these findings in simulated and real-data.
Proofs and additional experiments are in the appendix. These results hold for any conditional sampler.
@@ 2.1 Setting
We observe $\{(x_{i},y_{i})\}_{i=1,\ldots,n_{\mathrm{test}},\ldots,n_{\mathrm{test}}+n_{\mathrm{train}}}$ a test and train set sampled i.i.d. from $P_{0}$, a distribution from a class of distributions $\mathcal{M}$, where $X\in\mathcal{X}\subset\mathbb{R}^{p}$ and $y\in\mathcal{Y}\subset\mathbb{R}$. We are interested in assessing the importance of a coordinate $j\in\{1,\ldots,p\}$. We denote by $X^{j}$ the $j$-th column of $X$ and by $X^{-j}$ the vector $X$ with the $j$-th coordinate excluded. We consider a space of functions $\mathcal{F}$ with a norm $\|\cdot\|_{\mathcal{F}}$. The notation is summarized in the glossary (Appendix A).
Similarly to the Model-X framework (Barber and Candès, 2015; Candès et al., 2018), we want to avoid making strong assumptions about the relationship between inputs and outputs, as it can be complex. However, we assume that the relationship among the input covariates is simple, typically because they originate from the same generative process (e.g. a measurement device). Then, it can be much more efficient and accurate to study the relationship between the $j$-th covariate, $X^{j}$, given the rest, $X^{-j}$, rather than directly regressing $y$ on $X^{-j}$. This approach is also advantageous in settings where labeling data is expensive, as there are typically more available samples for the input pair $(X^{-j},X^{j})$ than for the pair $(X^{-j},y)$. However, CPI’s empirical robustness to this assumption is shown in (Paillard et al., 2025).
We denote by $m(X):=\mathbb{E}[y\mid X]$ (resp. $m_{-j}(X^{-j}):=\mathbb{E}[y\mid X^{-j}]$) the conditional expectation of the input $y$ given $X$ (resp. $X^{-j}$), by $\widehat{m}$ (resp. $\widehat{m}_{-j}$) its estimation. Similarly, we denote by $\nu_{-j}(X^{-j}):=\mathbb{E}[X^{j}\mid X^{-j}]$ and $\widehat{\nu}_{-j}$ its estimate. We denote by $\ell:\widehat{\mathcal{Y}}\times\mathcal{Y}\to\mathbb{R}$ the loss used to compare a prediction $\widehat{m}(X)$ with the output $y$. These estimations are the usual objective of machine learning models. This is obvious for the quadratic loss, but for many other losses ($R^{2}$, deviance, classification accuracy, and the area under the ROC curve, see Williamson et al. (2023)) the minimizer is also a function of this conditional expectation.
@@
We define this quantity as the (Generalized) Total Sobol Index because, when the loss is the quadratic loss, it corresponds to the unnormalized TSI (Homma and Saltelli, 1996). It represents the loss decay when the covariate is not used. It has also been referred to as the Generalized ANOVA (Williamson et al., 2021) and is considered a standard importance measure (Lei et al., 2018; Rinaldo et al., 2019; Hooker et al., 2021; Bénard et al., 2022; Williamson et al., 2023). A large decay indicates that the feature is useful, whereas a small one indicates that it lacks predictive power when the remaining covariates are used. Bénard et al. (2022) considered it as the best quantity for support recovery.
@@ 2.2 Related work
Nonparametric theory (Kennedy, 2023) shows that a simple plug-in estimate does not achieve optimal convergence, typically requiring a one-step correction. However, Williamson et al. (2021) demonstrated that this correction is unnecessary with the quadratic loss, and Williamson et al. (2023) extended this to other losses under certain regularity conditions. This provides valid confidence intervals. Similar results will be established for our method in Section 4.1.
A major practical limitation is that it requires retraining a model, $\widehat{m}_{-j}$, for each coordinate $j$. This is computationally intensive and it also introduces optimization errors that do not compensate as desired, as discussed in Section 3.2.
A first naive permutation-based approach consists in comparing the performance of $\widehat{m}$ on a test set where the $j$-th column is shuffled: the formal definition of PFI is given in Appendix B. Even though it avoids refitting and Mi et al. (2021) report good performance, it suffers from extrapolation bias, as predictions may fall in low-density regions where $\widehat{m}$ was not trained (Hooker et al., 2021). It also fails to control type-I error with highly correlated covariates (Chamma et al., 2024a). Moreover, Bénard et al. (2022) showed that it does not target an interpretable quantity: it decomposes into three components, only the first being desirable ($\psi_{\mathrm{TSI}}$). The other are misleading with correlation.
To address this issue, Strobl et al. (2008) proposed conditionally permuting the $j$-th coordinate with respect to the others. In this way, only the information exclusively given by $X^{j}$ is shuffled, while the relationship with the rest of the coordinates is preserved. As a result, one can make predictions based on the new conditional sample, denoted by $\widetilde{x}^{\prime(j)}_{i}$, where the apostrophe indicates that it was estimated. More formally, Chamma et al. (2024a) proposed:
@@
CPI requires training a separate regressor $\widehat{\nu}_{-j}$ for each covariate, which may seem counterintuitive since permutation approaches aim to avoid fitting $\widehat{m}_{-j}$. However, as discussed in Section 2.1, while the relationship between $y$ and $X$ can be complex, between $X^{j}$ and $X^{-j}$ is simple. Then, unlike with LOCO, we use a simple model to estimate $\nu_{-j}$.
The goal of CPI is to sample $\widetilde{X}^{(j)}$ such that $\widetilde{X}^{(j)}\sim X,\widetilde{X}^{(j)-j}=X^{-j}$ and $\widetilde{X}^{(j)j}\perp\!\!\!\perp y\mid X^{-j}$. However, Chamma et al. (2024a) did not discuss the assumptions under which this method samples from the target distribution; we address this in Section 3.1. Furthermore, CPI does not estimate a known theoretical quantity, so in Section 4 we propose a modification to target TSI.
Finally, in Chamma et al. (2024a), a type-I error control is proposed using asymptotic normality. However, it does not address the fact that, under the conditional null hypothesis, similarly to what happens with LOCO, the influence function vanishes. Consequently, there is a need to correct the estimated variance to ensure type-I error control (Williamson et al., 2023; Dai et al., 2024; Verdinelli and Wasserman, 2024). In Section 5, we address this issue for valid inference.
@@ 3.1 Validity of conditional sampling
We observe that there is no need to preserve the same sampler used for CPI. Normalizing flows (Papamakarios et al., 2021) are theoretically sound, but impractical due to the large number of training samples they require. Other alternatives are domain-specific generative models, e.g. Sesia et al. (2020). Since the conditional sampling used in CPI, as well as in Hooker et al. (2021) and Blain et al. (2025), works well in practice and improves computational efficiency, we study its validity. Let us denote the $i$-th observation in which only the $j$-th coordinate has been conditionally permuted:
$\widetilde{x}^{{}^{\prime}(j)l}_{i}=\begin{cases}x^{l}_{i}&\text{if }l\neq j\\ \widehat{\nu}_{-j}(x_{i}^{-j})+\left\{x^{j}-\widehat{\nu}_{-j}(x^{-j})\right\}^{\mathrm{perm}}&\text{if }l=j\end{cases}$ (1)
@@
For each $j\in\{1,\ldots,p\}$, there exists a function $\nu_{-j}$ such that $X^{j}=\nu_{-j}(X^{-j})+\epsilon_{j}$ with $\epsilon_{j}\perp\!\!\!\perp X^{-j}$ and $\mathbb{E}\left[\epsilon_{j}\right]=0$
@@
For instance, the Random Forest is consistent under mild assumptions (Scornet et al. (2015)). If we assume that $X$ is Gaussian, a linear model is consistent. To deal with high-dimensionality, we generally use a Lasso (Tibshirani, 1996).
@@
Proposition 3.3 implies that CPI’s sampler is asymptotically drawing from the desired distribution. Then, CPI qualifies as a Conditional Feature Importance (Hooker et al., 2021). However, in the following, we abstract from this choice and provide general results for any empirical sampler $P^{\prime}_{j}$.